OpenStudy (anonymous):

did i get this right? The area of this cylinder's base is one hundred fifty square inches. The height of the cylinder is 6 inches. What is the volume of the cone that has the same base area and height? Show the steps of your solution and explain your work. cylinder having a height of 6 inches The pyramids at Giza include several square-based pyramids that were built in honor of the Egyptian gods. One of these incredible monuments is the Khafre pyramid, built as part of a temple complex, with the famous Sphinx at its base. The base length of this pyramid is seven hundred five feet, and it

OpenStudy (anonymous):

first question answer: Answer=5,828 Square Inches Steps: 1. 1/3 x PI x 75x75 2. 1/3 x 3.14 x 75x75 3 .33 x 3.14 x 5,625 4. answer= 5,828 I just used a formula I looked up on google, and that's how I got the answer. Plus I got help from my brother. second question answer:The answer is 110,685 Feet steps: BxH/3 705 x 471/3 110,685 ft i looked up the formula again on google.

OpenStudy (anonymous):

@some_someone

Directrix (directrix):

@nickmifsud1223 The volume of a cylinder is: V = B*h where B is the area of the base and h is the height of the cylinder So, the volume of the cylinder is 150*6 = ? ---------------- Because the volume of a cone is 1/3 * B * h and the cylinder and the cone have the same base area and same height, would the following be true: Volume of cone = 1/3* volume of cylinder ? If so, the cone's volume would be 1/3 of 150*6 which is not 5,828. Note that volume should be expressed in cubic units.